### v1

@@ 4.2.2 Impact of Missing Preprocessing Operations
(a) Numerical Instability: Models trained on raw text data showed clear signs of numerical instability. According to our analysis, 63.33% of the models exhibited this issue. From our analysis, we observed that 80-90% of their layers were affected, particularly in the embedding layers and convolutional layers. These layers showed extreme weight ranges (-2.16 to 2.54), significantly deviating from the typical ranges of -0.1 to 0.1 for weights observed in the models trained on the clean dataset. These extreme weight parameter values indicate potential instability in the models and can be traced to the lack of proper text preprocessing and tokenization brownlee_2017 . Without these steps, the model encounters a high frequency of out-of-vocabulary words, including typos, rare technical terms, and inconsistent word forms (e.g., different tenses or misspellings of the same word). As the model attempts to learn representations for these out-of-vocabulary words, it often resorts to extreme weight adjustments, particularly in the embedding and convolutional layers, which are responsible for feature extraction. This leads to high volatility and significant variance in the layer weights, which results in exploding gradients and numerical instability in the neural networks bengio1994learning ; goodfellow2016deep .
(b) Skewed Bias Distributions: Models trained on raw texts demonstrated significant skewness in their bias distributions. Our analysis revealed that 53.33% of the models exhibited this issue, with 50-55% of the layers affected, particularly the convolutional layers. These layers demonstrated high skewness ($>$ 1) and moderate kurtosis ($|k|>$ 3) in their bias values, which are significantly higher than their baseline counterparts (skewness = 0.34, kurtosis = 1.82), collected from the models trained on clean, preprocessed data. This phenomenon can be attributed to the imbalanced representation of the text data due to a lack of preprocessing operations. Without tokenization and stop word removal, certain stop words or phrases may appear disproportionately in the dataset. This imbalance directly affects the bias values in the model. As certain words or phrases occur more frequently, the neurons associated with these common patterns are activated more often. Consequently, the biases for these neurons receive more frequent updates during training, pushing them further from their initial values. In contrast, biases for neurons associated with less common words or patterns receive fewer updates. This disparity in update frequency leads to a spread in the bias values, with some shifting significantly and others remaining closer to their starting points. This results in a skewed distribution of bias values, which reflects the imbalanced nature of the input data.
@@
4.2.3 Qualitative Analysis: Impact of Concept Drift on Decision Boundary
@@
To assess the impact of data quality issues on deep learning models’ training behaviour, we also performed a qualitative analysis using t-SNE plots. First, we trained 30 models on clean datasets and 30 models on datasets containing concept drift. We selected concept drift for our qualitative analysis, as it affected the highest number of models (see Table. 5). Then, we generated the t-SNE plots for the dense layer representations, as shown in Fig. 2 and Fig. 3. We focus on dense layers since their decision boundaries capture complex, non-linear relationships between features and directly contribute to the final classification. Furthermore, existing literature messaoud2022duplicate has analyzed the t-SNE plots of dense layers in their qualitative analysis. We discuss the outcome of our analysis below.
@@
From Fig. 2 and 3, we note the impact of concept drift on the dense layers of our analyzed deep learning models for duplicate bug report detection. In the first dense layer (Fig. 2(a)), models trained on clean data show a clear separation between the two classes, indicating effective capture of the patterns. On the other hand, models trained on buggy data (Fig. 3(a)) exhibit complex representations with ambiguous decision boundaries. This observation is supported by an analysis of KL Divergence, which serves as the cost function of the t-SNE plot. As the cost function, KL Divergence is automatically calculated and optimized during the generation of these visualizations, with lower KL Divergence values indicating a better representation of the high-dimensional data in the 2D space. For the first dense layer, the mean KL divergence for buggy models (0.32) is higher than that of bug-free models (0.26), indicating more noise in the layer representation of models affected by concept drift. The second dense layer (Fig.2(b) and Fig.3(b)) demonstrates similar trends, with bug-free models showing well-defined decision boundaries and buggy data models displaying unclear decision boundaries and fragmented clusters. This trend is further corroborated by the KL divergence analysis, where buggy models exhibit a mean KL divergence of 0.24 compared to 0.14 for bug-free models in the second dense layer. The increase in KL divergence for buggy models across both layers (23% in the first layer and 71% in the second layer) underscores the concept drift’s impact on the network. It also highlights how concept drift progressively affects a model’s ability to learn meaningful representations and make accurate predictions, ultimately impairing its performance in duplicate bug report detection.
@@
Summary of RQ2: Quality issues in text-based data introduce many challenges for deep learning models, including abnormal weight distribution, gradient skewness, and overfitting. Moreover, missing preprocessing techniques can lead to numerical instability and skewed bias distributions, worsening a models’ ability to learn effectively from the bug reports.
@@
Data Quality Issues Sparse Parameter Updates Vanishing Gradients Poor Optimization
@@ 4.3.1 Impact of Data Quality
(a) Sparse Parameter Updates: We observed sparse updates to the parameters of neural networks trained on low-quality metric data. As shown in Table 6, 76.67% of models trained with class imbalance exhibited this phenomenon. Our analysis revealed that 50-55% of their layers were affected, particularly the fully connected layers. In these layers, up to 85% of the neurons showed no change in their gradients. That is, 85% of the neurons in a particular layer have not received any updates and have effectively stopped learning. This phenomenon can be explained by the disproportionate representation of classes in imbalanced datasets. In such datasets, the model is exposed to significantly more examples from the majority classes, causing neurons associated with these classes to receive frequent and substantial gradient updates. Conversely, neurons detecting minority classes receive updates far less frequently and with smaller magnitudes. As a consequence, a significant portion of neurons, particularly those responsible for minority class features, become ‘dead’ and do not contribute to the learning process. Thus, the model struggles to learn representations for minority classes, leading to poor performance in these underrepresented categories.
(b) Vanishing Gradients: We also observed the problem of vanishing gradients in models trained on imbalanced datasets. As shown in Table 6, 83.33% of the models trained on imbalanced datasets exhibited this phenomenon. In these models, gradient magnitudes decreased exponentially from the output layer towards the input layer. Similarly, models trained on corrupted data experienced gradient drops by a factor of $10^{6}$ between the output and input layers. In contrast, models trained on clean data showed a moderate drop in gradients, with only a $10^{2}$ factor decrease from output to input. As discussed above, class imbalance leads to sparse parameter updates, which in turn contributes to the vanishing gradient problem. As neurons associated with minority classes receive infrequent and smaller updates, their gradient signals become progressively weaker as they propagate backwards from the output layer through the network. Consequently, the early layers of the network (e.g., input layers), crucial for learning fundamental features, receive negligible updates for minority class patterns and suffer from vanishing gradients. This phenomenon severely impairs a model’s ability to learn from minority classes, further exacerbating the issues caused by class imbalance.
(c) Poor Optimization: We also noticed poor optimization in the models trained on low-quality metric data. As shown in Table 6, 66.67% of models trained on datasets with class imbalance suffered from this issue. Our analysis revealed that the models trained on low-quality data led to a stagnation of loss values after the first epoch. This contrasts sharply with models trained on balanced datasets, which exhibited steady and consistent decreases in loss throughout training. The average training and validation loss for models trained on imbalanced datasets was approximately 1.5 times higher than those trained on clean datasets. This trend was consistent across different software projects. For instance, in the Openstack project, the training loss value increased from $\approx$75 for clean data to $\approx$115 for buggy data. Similarly, in the QT project, the training loss value rose from $\approx$110 for clean data to $\approx$170 for buggy data. When a dataset has a class imbalance, the model’s corresponding optimization process is severely affected. The stagnation in loss values, coupled with the vanishing gradient problem, suggests that the model might get stuck in local minima. As a result, the model struggles to learn representations for minority classes or refine its decision boundaries, leading to poor performance.
@@ 4.3.2 Impact of Missing Preprocessing Operations
(a) Exploding Gradients: Models trained on raw numerical data showed signs of exploding gradients. According to our analysis, 66.67% of the models suffered from this issue. We observed that 70-80% of their layers were affected, particularly in the output layers of the network. These layers showed extreme gradient values (e.g., from -4.23 to 2.76), significantly deviating from the typical gradient ranges of -1 to 1 observed in the models trained on properly scaled datasets. These extreme gradient values can be explained by the lack of appropriate feature scaling and normalization glorot2010understanding . Without these preprocessing steps, the model encounters features with vastly different scales, leading to disproportionate weight updates. This leads to high volatility and significant variance in the gradient magnitudes, which results in exploding gradients and potential divergence from the normal behaviours in the neural networks pascanu2013difficulty ; goodfellow2016deep .
(b) High Variance in Weight Distribution: Models trained on unprocessed numerical data demonstrated significant variance in weight distribution. Our analysis revealed that 53.33% of the models exhibited this issue, with 85-90% of their layers affected, particularly the input and early hidden layers. These layers demonstrated high variance in weight magnitudes ($\sigma^{2}\approx 5.76$), which is significantly higher than their baseline counterparts ($\sigma^{2}\approx 0.89$) collected from the models trained on normalized data. This high variance in weight distribution stems from the interaction within a neural network and the diverse scales inherent in software metrics. For example, lines of code might be in thousands, while cyclomatic complexity might be in single digits. During training, the neural network compensates for these scale disparities to ensure a fair representation of all features in the learning process lecun2002efficient . It assigns larger weights to small-scale features (e.g., cyclomatic complexity) to amplify their effect and smaller weights to large-scale features (e.g., lines of code) to prevent them from dominating the output. This compensation mechanism results in a wide range of weight values. The input and early hidden layers are particularly susceptible to this issue as they directly interact with the raw feature values lecun2002efficient , contributing to the observed weight variance issue.
@@
4.3.3 Qualitative Analysis: Impact of Class Imbalance on Feature Importance
To analyze the impact of class imbalance on the buggy and bug-free models, we apply the GradCAM (Gradient-weighted Class Activation Mapping) technique selvaraju2017grad . This approach has been used in previous studies to understand the effectiveness and efficacy of defect prediction models issta2021deepjit . For our GradCAM analysis, we used datasets specifically designed for defect prediction, focusing on software metrics. Our dataset exhibited an imbalanced distribution, with the minor class comprising non-defective instances and the major class consisting of defective instances. By examining the GradCAM outputs, we investigate which metrics the models prioritize and how this focus differs between the two types of models trained on two distinct codebases: Openstack and QT.
@@ 4.3.3 Qualitative Analysis: Impact of Class Imbalance on Feature Importance
Buggy Models: Our analysis of the GradCAM output for the buggy models reveals a disproportionate focus on obscure or overly specific tokens across both codebases. In the OpenStack model, tokens like ‘audit_location_generator’ (0.9987) and ‘i62ce43a330d7ae94eda4’ (0.9962) have probabilities very close to 1, while more general programming concepts like ‘if’ (0.0023) and ‘return’ (0.0041) have probabilities near 0. Similarly, the Qt model exhibits high probabilities for tokens such as ‘Q_OBJECT_FAKE’ (0.9976) and ‘qAsConst’ (0.9953), while potentially bug-indicative tokens like ‘nullptr’ (0.0037) or ‘delete’ (0.0052) have very low probabilities. This pattern suggests that the severe class imbalance has likely led to overfitting, causing the models to associate rare, noise tokens, or codebase-specific terms with software defects in the training data. Consequently, these buggy models may fail to generalize and capture the broader semantics and context of software defects across different codebases.
@@
Bug-free Models: In contrast, according to GradCAM output, the bug-free models focus on a wider variety of relevant technical terms and phrases in both OpenStack and Qt codebases. For OpenStack, tokens such as ‘vendor-data’ (0.9871), ‘xenapinfs-glance-integration’ (0.9923), ‘live_migr-ate’ (0.9905), ‘deprecationwarning’ (0.9889), ‘regressions’ (0.9917), and ‘backward’ (0.9894) have probabilities close to 1. These terms intuitively relate to areas where bugs occurred in OpenStack code, such as migrations, deprecations, regressions, and vendor integrations. Similarly, for Qt, we observe a focus on both framework-specific and general programming concepts. Tokens like ‘QObject’ (0.9934), ‘connect’ (0.9912), ‘emit’ (0.9901), ‘virtual’ (0.9887), and ‘override’ (0.9923) have probabilities close to 1. Additionally, memory-related terms such as ‘new’ (0.9876), ‘delete’ (0.9908), and ‘shared_ptr’ (0.9892) also show high probabilities, reflecting the importance of memory management in C++ code. This observation suggests that the bug-free models learn to focus on both codebase-specific concepts and general programming patterns that might have historically been associated with software defects. By mitigating the data quality issues, particularly class imbalance, these models have developed more generalizable representations that meaningfully capture the semantics of software defects across different codebases and frameworks.
The stark contrast in token focus between the buggy and bug-free models highlights the significant impact of class imbalance on model behavior. While the buggy models emphasize seemingly arbitrary or overly specific tokens to learn about software defects, the bug-free models focus on relevant technical concepts and terms.
@@
Summary of RQ3: The presence of data quality issues in metric-based data introduces several challenges for deep learning models, including sparse parameter updates, vanishing gradients, and poor optimization. Moreover, inadequate preprocessing techniques can lead to exploding gradients and high variance in weight distribution, further hindering the models’ ability to learn from the data effectively.
@@
In code-based datasets, we found that 83.33% of models (25 out of 30) exhibited inconsistent learning, 86.67% (26 out of 30) showed reduced model capacity, and 76.67% (23 out of 30) experienced gradient instability when trained on data with label noise or concept drift. For text-based datasets, 93.33% of models (28 out of 30) showed abnormal weight distribution, 76.67% (23 out of 30) demonstrated gradient skewness, and 86.67% (26 out of 30) exhibited overfitting when trained on data with concept drift or class imbalance. In metric-based datasets, 76.67% of models (23 out of 30) showed sparse parameter updates, 83.33% (26 out of 30) experienced vanishing gradients, and 73.33% (22 out of 30) demonstrated poor optimization when trained on data with label noise or class imbalance. When models were trained without appropriate preprocessing of data, we also observed similar patterns across all dataset types. In code-based datasets, 66.67% of models (20 out of 30) trained without proper preprocessing showed slow convergence, and 56.67% (17 out of 30) exhibited incorrect learning. For text-based data, 66.67% of models (20 out of 30) trained on unprocessed data showed numerical instability, while 56.67% (17 out of 30) exhibited skewed bias distributions. In metric-based datasets, 73.33% of models (22 out of 30) trained without preprocessing demonstrated exploding gradients, and 56.67% (17 out of 30) showed high variance in weight distribution. All percentages observed in these new datasets align closely with our original findings, varying within a 1-5% margin. These consistencies in observations across diverse datasets underscore the generalizability of our findings, highlighting the persistent impact of data quality issues and the importance of proper preprocessing in model training for software engineering tasks.

### v2

@@ 4.2.1 Impact of Data Quality
CodeBERT + MLP
Label Noise 10.00% 40.00% 3.33%
Class Imbalance 43.33% 16.67% 83.33%
Concept Drift 70.00% 60.00% 33.33%
Table 9: Impact of Data Bug Symptoms on Model’s Performance for Text-Based Data
Symptoms Accuracy Precision Recall F1-Score
DCCNN
Baseline Performance 93.41% 91.68% 92.45% 92.11%
Abnormal Weight Distribution 88.23% 86.94% 85.17% 86.04%
Gradient Skewness 85.92% 84.13% 81.26% 82.67%
Overfitting 84.15% 82.31% 85.87% 84.06%
CodeBERT + MLP
Baseline Performance 89.62% 82.23% 88.97% 85.14%
Abnormal Weight Distribution 86.12% 79.23% 84.87% 81.94%
Gradient Skewness 83.96% 77.14% 82.31% 79.63%
Overfitting 82.17% 74.28% 85.92% 79.64%
@@ 4.2.2 Impact of Missing Preprocessing Operations
(a) Extreme Weights: When we trained our models without appropriate preprocessing operations, we observed extreme weight values across multiple layers. As shown in Table 10, 66.67% of DCCNN and 73.33% of CodeBERT+MLP models trained without stop word removal demonstrated extreme weights. Our manual analysis shows that 80-90% of their layers were affected, especially embedding and convolutional layers. The models trained without appropriate preprocessing operations exhibited extreme weight ranges (e.g., -2.16 to 2.54), significantly deviating from the stable values in baseline models (e.g., -0.1 to 0.1). These extreme weights also had a notable impact on the performance of the models. As shown in Table 11, DCCNN witnessed a 9.18%-9.47% decline across various performance metrics. CodeBERT+MLP exhibited similar behavior, with performance metrics metrics decreasing by 8.38%-8.74%.
(b) Skewed Bias Distributions: We also observed significant skewness in the bias distribution of our models trained without appropriate preprocessing. As shown in Table 10, 53.33% of DCCNN and 60.00% of CodeBERT+MLP models trained without stop word removal demonstrated skewed bias distributions. Our analysis of these models revealed that 50-55% of their layers were affected, particularly the convolutional layers. These layers had highly skewed bias values (e.g., $>1$) with moderate kurtosis (e.g., $|k|>3$), significantly different from the balanced distributions observed in models trained on properly preprocessed data (skewness: 0.34, kurtosis: 1.82). These skewed bias distributions also had a substantial impact on the performance of the models, causing 6.49%-7.17% and 5.75%-6.47% performance drops in various metrics (DCCNN, CodeBERT+MLP).
Table 10: Manifestations of Missing Preprocessing in Text-Based Models
Missing Preprocessing Operation Extreme Weights Skewed Bias
DCCNN
Stemming 63.33% 50.00%
Stop Word Removal 66.67% 53.33%
Case Conversion 60.00% 46.67%
CodeBERT + MLP
Stemming 70.00% 56.67%
Stop Word Removal 73.33% 60.00%
Case Conversion 66.67% 53.33%
Table 11: Impact of Missing Preprocessing on Model’s Performance for Text-Based Data
Symptoms Accuracy Precision Recall F1-Score
DCCNN
Performance on Clean Data 93.41% 91.68% 92.45% 92.11%
Extreme Weights 84.23% 82.14% 83.98% 83.05%
Skewed Bias Distributions 86.92% 84.73% 85.16% 84.94%
CodeBERT + MLP
Performance on Clean Data 89.62% 82.23% 88.97% 85.14%
Extreme Weights 81.24% 75.16% 80.23% 77.61%
Skewed Bias Distributions 83.87% 76.92% 82.14% 79.44%
@@
4.2.3 Qualitative Analysis: Impact of Concept Drift in Training Data on Decision Boundary
@@
To assess how the quality issues in text data affect deep learning models’ training behaviour, we also performed a qualitative analysis using t-SNE plots. First, we trained 30 models on clean datasets and 30 models on datasets containing concept drift. We selected concept drift for our qualitative analysis, as it affected the majority of models (see Table. 8). Then, we generated the t-SNE plots for the dense layer weights, as shown in Fig. 3 and Fig. 4. We focus on dense layers since their decision boundaries capture complex, non-linear relationships between features and directly contribute to the final classification. Furthermore, existing literature messaoud2022duplicate has used the t-SNE plots of dense layers in their qualitative analysis. We discuss the outcome of our analysis below.
@@
From Fig. 3 and 4, we note the impact of concept drift on the dense layers of our analyzed deep learning models for duplicate bug report detection. In the first dense layer (Fig. 3(a)), models trained on clean data show a clear separation between the two classes (duplicate and non-duplicate), indicating effective capture of the patterns. On the other hand, models trained on buggy data (Fig. 4(a)) exhibit complex feature representations with ambiguous decision boundaries. This observation is supported by an analysis of KL Divergence, which serves as the cost function of the t-SNE plot. KL Divergence is automatically calculated and optimized during the generation of these visualizations, with lower KL Divergence values indicating a better representation of the high-dimensional data in the 2D space. For the first dense layer, the mean KL divergence for bug-free models (0.26) is lower than that of buggy models (0.32), indicating more noise in the layer representation of models affected by concept drift. The second dense layer (Fig. 3(b) and Fig. 4(b)) demonstrates similar trends, with bug-free models showing well-defined decision boundaries and buggy data models displaying ambiguous decision boundaries and fragmented clusters. This trend is further corroborated by the KL divergence analysis, where bug-free models exhibit a mean KL divergence of 0.14 compared to 0.24 for buggy models in the second dense layer. The increase in KL divergence for buggy models across both layers demonstrates the concept drift’s impact on the network. It also highlights how concept drift progressively affects a model’s ability to learn meaningful representations and make accurate predictions, ultimately impairing its performance in duplicate bug report detection.
@@
Summary of RQ2: Quality issues in text-based data introduce many challenges for deep learning models, including abnormal weight distribution, gradient skewness, and overfitting. Moreover, missing preprocessing techniques can lead to extreme weights and skewed bias distributions, worsening the model’s ability to learn effectively from the bug reports.
@@
Data Quality Issues Sparse Parameter Updates Vanishing Gradients Higher Loss
DeepJIT
@@
CodeBERTJIT
Label Noise 53.33% 56.67% 40.00%
Class Imbalance 83.33% 90.00% 73.33%
Concept Drift 23.33% 26.67% 16.67%
Table 13: Impact of Data Bug Symptoms on Model’s Performance for Metric-Based Data
Symptoms Accuracy Precision Recall F1-Score
DeepJIT
Baseline Performance 73.14% 19.36% 71.02% 31.26%
Sparse Parameter Updates 69.23% 18.14% 67.86% 28.37%
Vanishing Gradients 64.12% 16.28% 60.34% 25.46%
Higher Loss 65.87% 17.23% 61.92% 26.18%
CodeBERTJIT
Baseline Performance 89.59% 36.78% 56.71% 39.45%
Sparse Parameter Updates 84.87% 34.23% 52.16% 35.28%
Vanishing Gradients 80.24% 29.15% 46.32% 35.47%
Higher Loss 81.93% 31.26% 48.75% 37.84%
@@ 4.3.1 Impact of Data Quality
(a) Sparse Parameter Updates: When we trained our models using buggy metric data, we observed sparse parameter updates in several layers. Normal learning is characterized by active neurons with regular gradient updates, whereas sparse updates to model parameters (e.g., gradients, weights, biases) suggest learning difficulties li2024neurrev . As shown in Table 12, 80.00% of the models trained on data with class imbalance demonstrated sparse parameter updates. Our manual analysis shows that 50-55% of their layers were affected, especially the fully connected dense layers, where a significant number of neurons remained inactive due to data imbalance. In contrast, models trained on balanced data enjoyed regular updates to their neurons and demonstrated consistent learning behavior. We also observed performance deterioration in the models with sparse updates. As shown in Table 3, DeepJIT witnessed 4.45-9.24% performance drop across various metrics. CodeBERTJIT exhibited similar behaviour with a decrease of 5.27-10.57% across various performance metrics.
(b) Vanishing Gradients: We observed severe gradient vanishing problem in our models trained on imbalanced datasets. According to existing studies, smaller gradients hinder information flow within a model and introduce learning difficulties pascanu2013difficulty ; goodfellow2016deep . Vanishing gradients occur when the gradient signal becomes progressively smaller as it propagates backward through the network, preventing effective weight updates in earlier layers. This phenomenon limits the network’s ability to learn long-range dependencies and complex patterns in the data. As shown in Table 12, 86.67% of models trained on imbalanced data demonstrated vanishing gradients. Our analysis of these models revealed that gradients decreased exponentially across the layers, with input layers receiving gradients $10^{6}$ times smaller than output layers. This reduction significantly differs from that observed in models trained on clean data (e.g., factor of $10^{2}$). These vanishing gradients had a significant impact on the model’s performance. As shown in Table 13, DeepJIT witnessed a performance reduction of up to 12.33-18.55% across the various performance metrics. CodeBERTJIT exhibited similar degradation with 10.09-20.74% decreases across all performance metrics.
(c) Higher Loss: When we trained the models using buggy metric data, we also observed significantly higher training losses. From the perspective of neural networks, optimal learning suggests decreasing loss values as the models progressively fit the training data goodfellow2016deep . Higher persistent loss indicates a model’s struggle to find optimal parameters that minimize the objective function. As shown in Table 12, 70.00% of models trained with class imbalance demonstrated this issue. These models showed stagnant losses after the first epoch, with average losses $\approx$1.50 times higher than baseline models. For instance, training loss increased from $\approx$75 to $\approx$115 in OpenStack and from $\approx$110 to $\approx$170 in QT projects. This resulted in performance deterioration, DeepJIT witnessed 9.94-16.25% relative drop across performance metrics. CodeBERTJIT exhibited similar degradation with 4.08-15.01% relative decreases across all performance metrics (Table 13).
@@ 4.3.2 Impact of Missing Preprocessing Operations
(a) Exploding Gradients: When we trained our models without normalizing their features, we observed extreme gradient values across multiple layers. As shown in Table 14, 70.00% of DeepJIT and 76.67% of CodeBERTJIT models trained without feature normalization demonstrated exploding gradients. Our manual analysis shows that 70-80% of their layers were affected, especially output layers, which exhibited extreme gradient values (e.g., -4.23 to 2.76), significantly deviating from the stable values (e.g., -0.95 to 0.86) observed in baseline models. We also observed notable performance deterioration in models with exploding gradients. As shown in Table 15, DeepJIT witnessed a 9.87%-15.13% decline across various performance metrics. CodeBERTJIT exhibited similar behavior, with metrics decreasing by 5.55%-17.81%.
(b) High Variance in Weight Distribution: We observed significant variance in the weight distributions of our models trained without normalized features. According to existing studies, balanced parameter distributions typically show moderate variance ($\sigma^{2}\approx 0.89$lecun2002efficient . As shown in Table 14, 63.33% of DeepJIT and 70.00% of CodeBERTJIT models trained without feature normalization demonstrated high weight variance. Our analysis of these models revealed that 85-90% of their layers were affected, particularly input and early hidden layers. These layers had high variance values (e.g., $\sigma^{2}\approx 5.76$), significantly different from that of the models trained with normalized features. These high-variance weight distributions also had a noticeable impact on the model’s performance, causing a 7.81%-11.58% and 0.43%-11.77% drops in various performance metrics for our baseline techniques (DeepJIT, CodeBERTJIT).
Table 14: Impact of Missing Preprocessing in Metric-Based Models
Missing Operation Exploding Gradients Weight Variance
DeepJIT
Feature Normalization 70.00% 63.33%
Feature Scaling 66.67% 60.00%
CodeBERTJIT
Feature Normalization 76.67% 70.00%
Feature Scaling 73.33% 66.67%
Table 15: Impact of Missing Preprocessing on Model’s Performance for Metric-Based Data
Symptoms Accuracy Precision Recall F1-Score
DeepJIT
Performance on Clean Data 73.14% 19.36% 71.02% 31.26%
Exploding Gradients 65.92% 16.87% 63.14% 26.53%
High Variance in Weight Distribution 67.43% 17.56% 65.23% 27.64%
CodeBERTJIT
Performance on Clean Data 89.59% 36.78% 56.71% 39.45%
Exploding Gradients 82.14% 30.23% 48.92% 37.26%
High Variance in Weight Distribution 84.37% 32.45% 50.16% 39.28%
@@
4.3.3 Qualitative Analysis: Impact of Class Imbalance in the Training Data on Feature Importance
GradCAM visualization has been used in software engineering research to analyze the feature importance in defect prediction models selvaraju2017grad ; issta2021deepjit . In our study, we also analyze GradCAM outputs to understand the impact of data quality issues on the model’s training. Among four data quality issues, we found that class imbalance had the strongest impact on models trained with metric-based data (see Table 12). Thus, we conduct a qualitative analysis to better understand how class imbalance affects a model during training. Towards this, we trained 60 models in total. We divided these models into four groups of 15 models each: (a) models trained on a clean Openstack dataset, (b) models trained on a clean QT dataset, (c) models trained on an Openstack dataset with class imbalance, and (d) models trained on a QT dataset with class imbalance. We discuss our findings below.
@@ 4.3.3 Qualitative Analysis: Impact of Class Imbalance in the Training Data on Feature Importance
Buggy Models: Our analysis of the GradCAM outputs targeting the buggy models reveals a disproportionate focus on obscure or overly specific tokens by the models. In the models trained on OpenStack dataset, tokens like ‘audit_location_generator’ (0.9987) and ‘i62ce43a330d7ae94eda4’ (0.9962) have probabilities very close to 1, while more general programming concepts like ‘if’ (0.0023) and ‘return’ (0.0041) have probabilities near 0. These probabilities indicate the influence of the tokens on the final output. So, the high probabilities of relevant tokens indicate that the model uses the tokens relevant to the defects in making the final decision. Similarly, the models trained on the Qt dataset exhibit high probabilities for tokens such as ‘Q_OBJECT_FAKE’ (0.9976) and ‘qAsConst’ (0.9953), while potentially salient tokens like ‘nullptr’ (0.0037) or ‘delete’ (0.0052) have very low probabilities. These patterns suggests that the dominant presence of negative instances (a.k.a., severe class imbalance) has likely led to overfitting, causing the models to associate rare, noisy tokens, or codebase-specific terms with software defects in the training data. Consequently, these buggy models may fail to generalize and capture the broader semantics and context of software defects across different codebases.
(a) GradCAM outputs for high-relevance tokens demonstrating clear differentiation between bug-free and buggy models.
(b) GradCAM outputs for low-relevance tokens showing opposite trends between bug-free and buggy models.
Figure 5: Analysis of GradCAM activation patterns across token relevance. The bug-free model (green) shows high probabilities for semantically relevant tokens and low for irrelevant ones, while the buggy model (red) exhibits inverse behaviour, validating the model’s ability to identify bug-indicative features.
@@
Bug-free Models: In contrast, according to GradCAM output, the bug-free models focus on a wider variety of relevant technical terms and phrases in both OpenStack and Qt codebases. For OpenStack, tokens such as ‘vendor-data’ (0.9871), ‘xenapinfs-glance-integration’ (0.9923), ‘live_migr-ate’ (0.9905), ‘deprecationwarning’ (0.9889), ‘regressions’ (0.9917), and ‘backward’ (0.9894) have probabilities close to 1. These terms intuitively relate to areas where bugs occurred in OpenStack code, such as migrations, deprecations, regressions, and vendor integrations. Similarly, for Qt, we observe a strong focus on both framework-specific and general programming concepts. Tokens like ‘QObject’ (0.9934), ‘connect’ (0.9912), ‘emit’ (0.9901), ‘virtual’ (0.9887), and ‘override’ (0.9923) have probabilities close to 1. Additionally, memory-related terms such as ‘new’ (0.9876), ‘delete’ (0.9908), and ‘shared_ptr’ (0.9892) also show high probabilities, reflecting the importance of memory management in C++ code. These observations suggest that the bug-free models learn to focus on both codebase-specific concepts and general programming patterns that might have historically been associated with software defects. With balanced training data, they might have developed more generalizable representations that meaningfully capture the semantics of software defects across different codebases and frameworks.
The GradCAM visualization in Figure 5 shows a striking contrast in how bug-free and buggy models process code tokens. For semantically relevant tokens, the bug-free models demonstrate consistently high probabilities (e.g., 0.99 for ’deprecationwarning’, ’regressions’, ’QObject’, and ’virtual’), while the buggy models show low probabilities (0.14-0.18). Conversely, for non-important tokens, the pattern inverts dramatically - the buggy models assign maximum probability (1.00) to noise tokens like ’i62ce43a330d7ae94eda4’. Whereas, the bug-free models maintain appropriately low probabilities for the same tokens (0.14-0.18). Moreover, both models show similar low probabilities (0.17-0.18) for basic programming tokens, suggesting the bug-free models have a better understanding of defective and non-defective tokens. The probabilities across the salient and noisy tokens demonstrate that the bug-free models have learned to effectively discriminate between salient features and noise, while the buggy models are skewed toward overly specific or irrelevant tokens, limiting their generalization capability across different codebases.
@@
Summary of RQ3: The presence of data quality issues in metric-based data introduces several challenges for deep learning models, including sparse parameter updates, vanishing gradients, and higher loss. Moreover, inadequate preprocessing techniques can lead to exploding gradients and high variance in weight distribution, further hindering the models’ ability to learn from the data effectively.
@@
In code-based datasets, we found that $81.67\%$ of models exhibited near-zero biases, $85.00\%$ showed smaller weights, and $68.33\%$ experienced gradient instability when trained on data with label noise or concept drift. For text-based datasets, $85.00\%$ of models showed abnormal weight distribution, $73.33\%$ demonstrated gradient skewness, and $85.00\%$ exhibited overfitting when trained on data with concept drift or class imbalance. In metric-based datasets, $80.00\%$ of models showed sparse parameter updates, $86.67\%$ experienced vanishing gradients, and $70.00\%$ demonstrated poor optimization when trained on data with label noise or class imbalance.
When models were trained without appropriate preprocessing of data, we also observed similar patterns across all dataset types. In code-based datasets, $70.00\%$ of models trained without proper preprocessing showed slow convergence, and $56.67\%$ exhibited extreme bias values. For text-based data, $68.33\%$ of models trained on unprocessed data showed extreme weight values, while $55.00\%$ exhibited skewed bias distributions. In metric-based datasets, $70.00\%$ of models trained without preprocessing demonstrated exploding gradients, and $56.67\%$ showed high variance in weight distribution. All percentages observed in these new datasets align closely with our original findings, varying within a $1$-$5\%$ margin. These consistencies in observations across diverse datasets underscore the generalizability of our findings, highlighting the persistent impact of data quality issues and the importance of proper preprocessing in model training for software engineering tasks.